OpenStudy (anonymous):

Can someone please help me out ? You would be such a lifesaver you don't even understand .... Prove that each equation is an identity : 1) sin0 ( the 0 looks like acursive lower case o ) csc0 cos0 = cos0. 2) tan0 sin0 cos0 = sin^2 0. 3) cot0/cosc0 = cos0. 4) Csc0/sec0 = cot0

OpenStudy (lgbasallote):

2) tan0sin0cos0 = sin^2 0 (sin0/cos0)sin0cos0 = sin^2 0 cancel cos 0 sin^2 0 = sin^2 0

OpenStudy (lgbasallote):

1) sin0 csc0cos0 = cos0 sin0(1/sin0)cos 0 = cos0 cancel sin 0 cos 0 = cos 0

OpenStudy (campbell_st):

I think 1 is a typo csc = 1/sin so its cos/sin = cot (3) cos/sin / 1/sin = cos/sin * sin/1 = cos

OpenStudy (campbell_st):

(4) 1/sin / 1/cos = 1/sin * cos/1 = cot